OpenStudy (wesdg1978):

Write z1 and z2 in polar form and then find the product z1 z2 if: z1=2(sqrt3+i) z2=3(1+i sqrt3)

OpenStudy (anonymous):

First magnitude using pythagoras' theorem. In both cases it is sqrt( sqrt(3)^2 + 1^2) = sqrt(4) = 2 Now work out the angles. Both terms are positive so they're in the first quadrant (0 < angle < pi/2). sine(angle) = complex_part / magnitude sin(angle1) = 1/2 => angle1 = pi/6 (30 degrees) sin(angle2) = sqrt(3)/2 => angle2 = pi/3 (60 degrees) To multiply, you multiply the magnitudes and add the angles. To divide, you divide the magnitides and subtract the angles. To do 1/z1, do the same thing, remembering "1" is magnitude 1, angle 0 degrees. So for example z1/z2 magnitude = 2/2 = 1 angle = pi/6 - pi/3 = -pi/6 => z1/z2 = magnitude 1, angle -pi/6 which is (though you don't need to do this) sqrt(3)/2 - i/2 Check with normal division (again you don't need to do this): z1/z2 = (sqrt(3) + i) / (1 + i sqrt(3) = (sqrt(3) + i)(1 - i sqrt(3) / [(1 + i sqrt(3)(1 - i sqrt(3)] = (2sqrt(3) - 2i) / 4 = sqrt(3)/2 - i/2 Do z1 x z2 and 1/z1 in the same way.